Total Points Used 3646 The plates of a paraliel plate capacitor are 4 cm wide and 9 cm long. The plates are separated by a 2.0-mm-thick layer of paper (a) Calculate the capacitance of the device (in pr) using the dielectric constant of paper from this table (b) Any dielectric material other thanh, vacuum has a maximum electric field that can be generated in the dielectric m The dielectric strength for paper is reached at a value of 15 x 106 V/m. Calulate the maximum charge (in pC) that can be placed on the
Answer #1

b)

the potential difference between the plates is, V= E*d = 15*106*2*10-3

V=30000 V

for a capacitor,

Q=C*V = 55.8*10-12*30000

Q=1.674*10-6 C = 1.674  \muC